# Monotone Neural Barrier Certificates

### v1

@@ Appendix E Proof of Theorem 4
    where the first inequality holds by the monotonicity of $N$ and $f$ and the fact that $x\leq\overline{x}^{k}$, the second inequality holds by monotonicity of $N$ and $f$ and the fact that $u\leq\max_{j\in\mathcal{S}}\overline{u}^{j}$. This means that $N$ satisfies condition (4).
@@
Therefore, the monotone neural network $N:\mathcal{X}\to$ satisfies conditions (2), (3), and (4). The result then follows using Theorem (1). This completes the proof of part 1.
@@
Regarding part 2, our goal is to show that the monotone neural network $N:\mathcal{X}\to$ satisfying conditions (7), (8), and (10) is a control barrier certificate for $\mathfrak{S}$ with respect to $\mathcal{X}_{u}$. More specifically, our goal is to show that $N$ satisfies conditions (2), (3), and (5).
  1. 1.
@@
    (NN satisfies condition (2)) Let $x\in\mathcal{X}_{0}$. Since $\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $k\in\mathcal{I}$ such that $x\leq\overline{x}^{k}$. Since $N:\mathcal{X}\to$ satisfies condition (7), we have $N(\overline{x}^{k})\leq\gamma$. On the other hand, NN is a monotone neural network, thus $N(x)\leq N(\overline{x}^{k})\leq\gamma$. This shows that the monotone neural network $N$ satisfies condition (2).
  2. 2.
@@
    (NN satisfies condition (3)) Let $x\in\mathcal{X}_{u}$. Since $\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $j\in\mathcal{U}$ such that $\underline{x}^{j}\leq x$. Since $N$ satisfies the condition (8), we have $\eta\leq N(\underline{x}^{j})$. Using the monotonicity of the neural network NN, we get $\eta\leq N(\underline{x}^{j})\leq N(x)$. This means that the monotone neural network $N$ satisfies condition (3).
  3. 3.
@@
    (NN satisfies condition (5)) Let $x\in\mathcal{X}$ be such that $N(x)\leq\delta$. Since the family of hyper-rectangles $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ covers the state set $\mathcal{X}$, there exists $k\in\mathcal{P}$ such that $x\in[\underline{x}^{k},\overline{x}^{k}]$. Using the monotonicity of $N$, we get $N(\underline{x}^{k})\leq N(x)\leq\delta$. Since $N$ satisfies condition (10), we have $\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta$. Since both $N$ and $f$ are monotone functions, for every $u\in U$, we have $N(f(x,u))\leq N(f(\overline{x}^{k},u))$. By taking the infimum of both sides of this inequality, we obtain
    $\inf_{u\in U}N(f(x,u))\leq\inf_{u\in U}N(f(\overline{x}^{k},u))\leq\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta,$ (24)
@@
    where the second inequality holds using the fact that $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ is a hyper-rectangular cover for $U$. The inequality (24) implies that $N$ satisfies condition (5).
@@
Therefore, the monotone neural network $N:\mathcal{X}\to$ satisfies conditions (2), (3), and (5). The result then follows using Theorem (2). This completes the proof of part 2.
Appendix F Algorithm for Safety Verification
In this appendix, we present an algorithm, analogous to Algorithm (1), for constructing monotone neural robust barrier certificates.
Algorithm 2 Monotone Neural Robust Barrier Certificate
Input: Simulator access to $f$, sets $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, $\mathcal{X}=[\underline{x}_{\min},\overline{x}_{\max}]$, $e$ total number of epochs, $\hat{e}$ number of epochs before sampling, $m$ number of samples;
Output: Neural network $B$ (barrier certificate);
1:Over-approximate $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, and $\mathcal{X}$ using unions of hyper-rectangles as in (6)
2:Construct maximal initial points $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ and minimal unsafe points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$
3:Construct the hyper-rectangular cover $T=[\underline{x}_{u},\overline{x}_{\max}]$
4:Construct $H$ as the set of maximal points of $\mathcal{X}\setminus T$
5:Initialize monotone neural networks $B$ and set hyperparameters $\gamma<\delta=\eta$
6:for $e$ number of epochs do
7:  Minimize loss $\widehat{L}_{\mathrm{total}}=L_{1}+L_{2}+L_{4}$ as defined in (11)- (13), respectively.
8:  if Conditions (7)–(9) not satisfied for $\hat{e}$ epochs then
9:   $Q\leftarrow$ Sample $m$ points from $\{x\in\mathcal{X}\mid\gamma<B(x)<\eta\}$
10:   Update $T\leftarrow T\cup\bigcup_{x\in Q}[x,\overline{x}_{\max}]$
11:   Update $H$ by getting maximal points of $\mathcal{X}\setminus T$
12:   Set $\delta\leftarrow\min_{x\in H}N(x)$
13:  end if
14:  if all constraints are satisfied then
15:   return trained network $B$
16:  end if
17:end for
18:return Inconclusive (training failed to converge)
Appendix G Correctness and Sample Complexity of Algorithm 1
In this appendix, we provide additional details and proofs for Algorithm (1) presented in Section (3.3). We begin by presenting a systematic approach for computing the maximal elements of the set $\mathcal{X}\setminus T$.
G.1 Computing Maximal Points of $\mathcal{X}\setminus T$
[…]
$T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}].$ (25)
Then, we can characterize the set $\mathcal{X}\setminus T$ as follows:
$\mathcal{X}\setminus T=\left\{y\in\mathcal{X}\,\middle|\,\forall x\in Q\cup\{\underline{x}_{u}\},\,\exists i\in\{1,\ldots,n\}\text{ such that }y_{i}<x_{i}\right\}.$ (26)
That is, a point $y\in\mathcal{X}$ lies in the region $\mathcal{X}\setminus T$ if it is strictly smaller than every $x\in Q\cup\{\underline{x}_{u}\}$ in at least one coordinate.
To conservatively cover $\mathcal{X}\setminus T$, we construct a set $\mathcal{H}$ of maximal points defining hyper-rectangular regions. This is done through a structured procedure as follows:
Step 1: Incomparable base sampling. We begin by sampling a set $Q=\{y^{1},\ldots,y^{m}\}\subset\mathcal{X}\setminus T$ such that all points are pairwise incomparable. These serve as the base points for constructing $\mathcal{H}$.
Step 2: Maximal combinations. For each nonempty subset $S\subseteq\{y^{1},\ldots,y^{m}\}$, we compute the element-wise maximum
$x_{S}:=\max\{y\mid y\in S\},$
and collect the resulting points into a set $\mathcal{H}_{1}$. These represent the upper envelope of the base points.
Step 3: Anchor points. To ensure full coverage of the domain boundaries, we define a set of anchor points. For each coordinate $i\in\{1,\ldots,n\}$, define $z^{i}\in\mathcal{X}$ by
$(z^{i})_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ \min_{y\in Q}y_{i}&j=i\end{cases}$
We collect the resulting points in the set $\mathcal{H}_{2}$, for every $i\in\{1,\ldots,n\}$.
Step 4: Dominance filtering. We define the final set $\mathcal{H}$ as
$\mathcal{H}:=\text{clipmax}(\mathcal{H}_{1}\cup\mathcal{H}_{2}),$ (27)
where $\text{clipmax}(\cdot)$ denotes the subset of points not dominated by any other in the union. That is, we remove all $x\in\mathcal{H}_{1}\cup\mathcal{H}_{2}$ such that there exists $x^{\prime}\neq x$ with $x^{\prime}\leq x$.
We will need the following Lemma to show that $\mathcal{H}$ consists of maximal points of the set $\mathcal{X}\setminus T$.
Lemma 4 (Complement of a hyper-rectangle).
Let $a\in\mathcal{X}$ and $A=[a,\overline{x}_{\max}]$ be a hyper-rectangle. Then we have
$\mathcal{X}\setminus A\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],$
where $\mathcal{G}=\{a\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ a_{i}&j=i\end{cases}$
Proof.
Let $w\in\mathcal{X}\setminus A$. This means that $w\not\geq a$. As a result, there exists $i\in\{1,\ldots,n\}$ such that $w_{i}<a_{i}$. Now consider the point $v^{i}$, for every $j\neq i$, we have $v^{i}_{j}=(\overline{x}_{\max})_{j}\geq w_{j}$. Moreover, we have $v^{i}_{i}=a_{i}>w_{i}$. This means that $v^{i}>w$ and therefore $w\in[\underline{x}_{\min},v^{i}]$. ∎
Now we extend the previous result for two hyper-rectangles.
Lemma 5 (Complement of union of two hyper-rectangles).
Let $a,b\in\mathcal{X}$ and $A=[a,\overline{x}_{\max}],B=[b,\overline{x}_{\max}]$ be two hyper-rectangles. Then we have
$\mathcal{X}\setminus(A\cup B)\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],$
where $\mathcal{G}=\{x_{S}\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where $x_{S}$ is element-wise maximum of $a$ and $b$ and for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ \min\{a_{i},b_{i}\}&j=i\end{cases}$
Proof.
We know that $\mathcal{X}\setminus(A\cup B)=(\mathcal{X}\setminus A)\bigcap(\mathcal{X}\setminus B)$. By Lemma 4, we have
$\mathcal{X}\setminus A\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],\qquad\qquad\mathcal{X}\setminus B\subseteq\bigcup_{\widehat{y}\in\mathcal{\widehat{G}}}[\underline{x}_{\min},\widehat{y}$],
where $\mathcal{G}=\{a\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ a_{i}&j=i\end{cases}$
and $\mathcal{\widehat{G}}=\{b\}\cup\bigcup_{i=1}^{n}\{\widehat{v}^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $\widehat{v}^{i}$ is defined as follows:
$\widehat{v}^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ b_{i}&j=i\end{cases}$
Therefore, we have
[…]
$=\left([\underline{x}_{\min},a]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},v^{i}\}]\right)\bigcap\left([\underline{x}_{\min},b]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\widehat{v}^{i}\}]\right)$
$\subseteq\left([\underline{x}_{\min},x_{S}]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},v^{i}\}]\right)\bigcap\left([\underline{x}_{\min},x_{S}]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\widehat{v}^{i}\}]\right)$
$\subseteq[\underline{x}_{\min},x_{S}]\cup\left(\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\min\{v^{i},\widehat{v}^{i}\}]\right),$
where the first equality follows from Lemma 4, the third inequality follows from the interval inclusions $\underline{x}_{\min},a]\subseteq[\underline{x}_{\min},x_{S}]$ and $\underline{x}_{\min},b]\subseteq[\underline{x}_{\min},x_{S}]$, and the last inequality follows from the distributive property of unions and intersections. ∎
Lemma 5 can be extended to an arbitrary number of hyper-rectangles by applying mathematical induction.
Theorem 6 (Maximal Points of $\mathcal{X}\setminus T$).
Let $\mathcal{X}\subseteq\mathbb{R}^{n}$, and let $T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}]$ be a finite union of upper hyper-rectangles. Suppose the set $\mathcal{H}$ is constructed through the procedure in Steps 1–4, as described in (27). Then
$\mathcal{X}\setminus T\subseteq\bigcup_{y\in\mathcal{H}}[\underline{x}_{\min},y].$
Proof.
The results follows by applying Lemma 5 inductively. ∎
Theorem 7 (Cardinality of $\mathcal{H}$).
Let $\mathcal{X}\subseteq\mathbb{R}^{n}$, and let $T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}]$ be a finite union of upper hyper-rectangles. Suppose the set $\mathcal{H}$ is constructed through the procedure in Steps 1–4, as described in (27). Then, the cardinality of the set $\mathcal{H}$ is $\mathcal{O}(n)$.
Proof.
By the procedure in Steps 1–4, we have $\mathcal{H}\subseteq\mathcal{H}_{1}\cup\mathcal{H}_{2}$. This implies that $|\mathcal{H}|\leq|\mathcal{H}_{1}|+|\mathcal{H}_{2}|$. Based on Step 2, we have $|\mathcal{H}_{1}|\leq 2^{m}-1$. Moreover, in Step 3, we construct $n$ anchor points. This means that $|\mathcal{H}_{2}|=n$. As a result, the cardinality of $\mathcal{H}$ is at most $2^{m}+n-1$ and therefore $\mathcal{O}(n)$. ∎
G.2 Proof of Theorem 5
Proof.
We divide the proof into two parts: (i) proof of correctness, and (ii) proof of linear sample complexity.
  1. 1.
    (Correctness). Suppose that Algorithm 1 converges and returns a monotone neural control barrier certificate $N$ and a feedback controller $K$. By the termination condition of Algorithm 1, $N$ must satisfy conditions (7)–(8). We now show that $N$ and $K$ also satisfy condition (9). Since the set $H$ consists of the maximal elements of $\mathcal{X}\setminus T$, and $\delta=\min_{x\in H}N(x)$, it follows that for every $x\in T$, we have $N(x)\geq\delta$. Therefore, condition (9) holds trivially on the region $T$ for the monotone neural control barrier certificate $N$ and the feedback controller $K$. Furthermore, the algorithm ensures that condition (9) also holds on $\mathcal{X}\setminus T$. Thus, $N$ and $K$ satisfy condition (9) over the entire state space $\mathcal{X}$. The result then follows directly from Theorem 4.
  2. 2.
    (Linear sample complexity). The sample complexity of Algorithm (1) is exactly the number of elements in the family of hyper-rectangles $\{[\underline{x}_{\min},\overline{x}^{i}]_{i\in\mathcal{P}}\}$ where $\overline{x}^{i}$ are the maximal points of the set $\mathcal{X}\setminus T$. By Theorem 7, the cardinality of the maximal points of $\mathcal{X}\setminus T$ is $|\mathcal{H}|$ and it is of $\mathcal{O}(n)$. This completes the proof of linear sample complexity.
Appendix H Urban Traffic Control
We consider an urban traffic network described by a family of links (roads) categorized according to their actuation type: $\mathcal{L}_{r}$ (roads with traffic light control), $\mathcal{L}_{o}$ (on-ramps controlled by ramp meters), and $\mathcal{L}_{f}$ (freeway segments without direct control). Uncontrolled roads are also treated like freeways, so the set of all links (roads) in this network is given by $\mathcal{L}=\mathcal{L}_{r}\cup\mathcal{L}_{o}\cup\mathcal{L}_{f}$. The number of vehicles on link $l$ at time $t$ is $x_{[l],t}\in[0,c_{l}]$, where $c_{l}$ is the capacity of link $l$. We assume that the traffic flow follows fluid-like behavior. Here, the state space is $\mathcal{X}:=\Pi_{l\in\mathcal{L}}[0,c_{l}]$. The outflow from link $l$ at time $t$ is:
$q_{[l],t}=\min\Big{\{}x_{[l],t},\;\bar{q}_{l},\;\min_{l^{\prime}:\tau(l^{\prime})=\eta(l)}\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}],t})\Big{\}},$ (28)
where $\bar{q}_{l}$ is max outflow from $l$, $\alpha_{l:l^{\prime}}$ is capacity ratio of link $l^{\prime}$ relative to $l$, and $\beta_{l:l^{\prime}}$ is turning ratio from $l$ to $l^{\prime}$. Figure 2 presents a schematic diagram of the link outflow $q[l]$ as a function of the state variables $x[l]$ and $x[l^{\prime}]$, commonly referred to as the fundamental diagram in traffic flow literature [23]. The flow $q[l]$ decreases when one or more downstream links lack sufficient capacity to accommodate incoming vehicles. In such cases—when the last term in (28) determines the minimum—we say the traffic is congested. Otherwise, if the upstream demand governs the flow, the traffic is considered to be in the free-flow regime. We define the congestion-free set $\Pi$ as:
$\Pi:=\left\{x\in\mathcal{X}\mid\min\{x_{[l]},\bar{q}_{l}\}\leq\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}]})\;\forall l,l^{\prime}:\tau(l^{\prime})=\eta(l)\right\}.$ (29)
The actual allowable outflow from link $l$ depends on its actuation:
$\tilde{q}_{[l],t}=\begin{cases}s_{[l],t}\cdot q_{[l],t}&\text{if }l\in\mathcal{L}_{r},\\ \min\{q_{[l],t},r_{[l],t}\}&\text{if }l\in\mathcal{L}_{o},\\ q_{[l],t}&\text{if }l\in\mathcal{L}_{f}.\end{cases}$ (30)
Here, $s_{[l],t}\in\{0,1\}$ indicates the traffic light status (1 for green, 0 for red), and $r_{[l],t}\in\mathbb{R}_{+}$ is the ramp metering rate, both of which are control input to the system. Thus, $U=\mathbb{R}_{>0}^{4}\times\{0,1\}^{14}$. To prevent conflicting greens, for any pair $l,l^{\prime}$ with perpendicular intersections, we require $s_{[l],t}+s_{[l^{\prime}],t}\leq 1$. The state transition map $f$ for this urban traffic network is defined by:
$x_{[l],t+1}=x_{[l],t}-\tilde{q}_{[l],t}+\sum_{l^{\prime}:\tau(l)=\eta(l^{\prime})}\beta_{l^{\prime}:l}\tilde{q}_{[l^{\prime}],t}:=f_{[l]}(x_{t},s_{t},r_{t}),\qquad\mbox{ for all }l\in\mathcal{L}.$ (31)
We define the set of unsafe states as $\mathcal{X}_{u}=\mathcal{X}\setminus\Pi$; that is, a state is considered unsafe if at least one link in the network is congested. Some trajectories of the system are depicted in Figure 3. We refer the reader to [42] for more details on model of this system, its topology, and proof of monotonicty. Parameters of the model are found in Table 3
$x_{[l]},x_{[l^{\prime}]}$$q_{[l]}$$\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}]})$free flowcongested flow
Figure 2: Fundamental diagram. The outflow from link $l$ decreases when the number of vehicles on its immediate downstream link $l^{\prime}$ approaches the downstream capacity. This reduction in flow, caused by limited downstream supply, characterizes the onset of congestion due to blocking behavior.
Figure 3: Trajectories of the system. Dotted lines indicate entering congestion status.
Table 3: Parameters of the urban traffic network
links parameters
1–26, 49–53 $\bar{q}=15$, $c=40$
27–34 $\bar{q}=15$, $c=30$
35–48 $\bar{q}=40$, $c=60$
Turning ratios value
$\beta_{2:50},\beta_{4,53},\beta_{8:51},\beta_{12:7},\beta_{13:28},\beta_{15:30},\beta_{16:28}$ 0.2
$\beta_{21:32},\beta_{24:32},\beta_{26:2},\beta_{36:31},\beta_{36:33},\beta_{39:27},\beta_{43:29}$
$\beta_{5:12},\beta_{6:13},\beta_{6:18},\beta_{10:21},\beta_{10:26}$ 0.3
$\beta_{1:20},\beta_{6:7}$ 0.4
$\beta_{1:2},\beta_{11:12},\beta_{14:30},\beta_{17:7},\beta_{17:18},\beta_{19:2},\beta_{19:20}$
$\beta_{31:22},\beta_{31:25},\beta_{33:24},\beta_{49:3},\beta_{49:50},\beta_{51:4},\beta_{52:5}$ 0.5
$\beta_{22:34},\beta_{23:24},\beta_{23:34},\beta_{27:14},\beta_{27:17},\beta_{29:16},\beta_{52:53}$
$\beta_{2:3},\beta_{3:4},\beta_{4:5},\beta_{8:9},\beta_{12:13},\beta_{13:14},\beta_{15:16},\beta_{16:17}$
$\beta_{20:21},\beta_{21:22},\beta_{24:25},\beta_{25:26},\beta_{36:37},\beta_{39:40}$ 0.8
$\beta_{43:44},\beta_{46:47}$
Capacity ratios value
$\alpha_{19:2},\alpha_{26:2},\alpha_{17:7},\alpha_{12:7},\alpha_{13:28},\alpha_{16:28}$ $0.5$
$\alpha_{14:30},\alpha_{15:30},\alpha_{21:32},\alpha_{24:32},\alpha_{22:34},\alpha_{23:34}$
Appendix I Freeway Traffic Verification
For the freeway traffic network $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ shown in Figure (4), the state space is given by $\mathcal{X}=[0,50]^{n}$, and the state transition map $f:\mathcal{X}\times U\to\mathcal{X}$ is defined by:
$f_{1}(q)$ $=q_{1}+\tau\left(F_{0\rightarrow 1}-F_{1\rightarrow 2}(q)\right),$
$f_{i}(q)$ $=q_{i}+\tau\left(F_{i-1\rightarrow i}(q)-F_{i\rightarrow i+1}(q)\right),\quad i\in\{2,\ldots,n-1\},$
$f_{n}(q)$ $=q_{n}+\tau\left(F_{n-1\rightarrow n}(q)-F_{n\rightarrow 0}(q)\right),$ (32)
where the input is the on-ramp inflow $F_{0\rightarrow 1}\in U=[0,9]$, the parameter $\tau=0.1$ is sampling time, and $q=\begin{bmatrix}q_{1}&\ldots&q_{n}\end{bmatrix}^{\top}$ with $q_{i}\geq 0$ representing the vehicle density in road segment $i$. The flow from road segment $i$ to $j$, denoted $F_{i\rightarrow j}(q)$, is defined as:
$F_{i\rightarrow j}(q)$ $=\min\left\{\varphi_{i}(q_{i}),\sigma_{j}(q_{j})\right\},$ (33)
$F_{n\rightarrow 0}(q)$ $=\varphi_{n}(q_{n}),$ (34)
where $\varphi_{i}(q_{i})$ models the increasing outflow demand of segment $i$, and $\sigma_{j}(q_{j})$ models the decreasing inflow supply of segment $j$. The demand and supply functions are given by:
$\varphi_{i}(q_{i})$ $=15\left(1-e^{-0.038q_{i}}\right),\quad\sigma_{i}(q_{i})=\mathrm{ReLU}\left(15\left(1-\left(\frac{q_{i}}{100}\right)^{4}\right)\right),\quad\forall i\neq 2,$
$\varphi_{2}(q_{2})$ $=10\left(1-e^{-0.138q_{2}}\right),\quad\sigma_{2}(q_{2})=\mathrm{ReLU}\left(10\left(1-\left(\frac{q_{2}}{50}\right)^{4}\right)\right).$ (35)
The initial set for this system is defined as $\mathcal{X}_{0}=[0,5]\times[0,18]\times[0,5]^{n-2}$. The unsafe set is $\mathcal{X}_{u}=[30,50]^{n}$, which indicates congestion in all roads. We refer the reader to [9] for more details on this system and proof of monotonicity. For $n=1000$, our Algorithm (F) find a monotone neural robust barrier certificate for $\mathfrak{S}$ in 135s (averaged over 30 runs).
Figure 4: A line digraph representing an $n$-dim traffic network.

### v3

@@ 4. Conclusions
    where the first inequality holds by the monotonicity of $N$ and $f$ and the fact that $x\leq\overline{x}^{k}$, the second inequality holds by monotonicity of $N$ and $f$ and the fact that $u\leq\max_{j\in\mathcal{S}}\overline{u}^{j}$. This means that $N$ satisfies condition (2.4).
@@
Therefore, the monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfies conditions (2.2), (2.3), and (2.4). The result then follows using Theorem (2.3). This completes the proof of part 1.
@@
Regarding part 2, our goal is to show that the monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying conditions (3.2), (3.3), and (3.5) is a control barrier certificate for $\mathfrak{S}$ with respect to $\mathcal{X}_{u}$. More specifically, our goal is to show that $N$ satisfies conditions (2.2), (2.3), and (2.5).
  1. (1)
@@
    (NN satisfies condition (2.2)) Let $x\in\mathcal{X}_{0}$. Since $\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $k\in\mathcal{I}$ such that $x\leq\overline{x}^{k}$. Since $N:\mathcal{X}\to\mathbb{R}$ satisfies condition (3.2), we have $N(\overline{x}^{k})\leq\gamma$. On the other hand, NN is a monotone neural network, thus $N(x)\leq N(\overline{x}^{k})\leq\gamma$. This shows that the monotone neural network $N$ satisfies condition (2.2).
  2. (2)
@@
    (NN satisfies condition (2.3)) Let $x\in\mathcal{X}_{u}$. Since $\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $j\in\mathcal{U}$ such that $\underline{x}^{j}\leq x$. Since $N$ satisfies the condition (3.3), we have $\eta\leq N(\underline{x}^{j})$. Using the monotonicity of the neural network NN, we get $\eta\leq N(\underline{x}^{j})\leq N(x)$. This means that the monotone neural network $N$ satisfies condition (2.3).
  3. (3)
@@
    (NN satisfies condition (2.5)) Let $x\in\mathcal{X}$ be such that $N(x)\leq\delta$. Since the family of hyper-rectangles $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ covers the state set $\mathcal{X}$, there exists $k\in\mathcal{P}$ such that $x\in[\underline{x}^{k},\overline{x}^{k}]$. Using the monotonicity of $N$, we get $N(\underline{x}^{k})\leq N(x)\leq\delta$. Since $N$ satisfies condition (3.5), we have $\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta$. Since both $N$ and $f$ are monotone functions, for every $u\in U$, we have $N(f(x,u))\leq N(f(\overline{x}^{k},u))$. By taking the infimum of both sides of this inequality, we obtain
    (B.1) $\inf_{u\in U}N(f(x,u))\leq\inf_{u\in U}N(f(\overline{x}^{k},u))\leq\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta,$
@@
    where the second inequality holds using the fact that $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ is a hyper-rectangular cover for $U$. The inequality (B.1) implies that $N$ satisfies condition (2.5).
@@
Therefore, the monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfies conditions (2.2), (2.3), and (2.5). The result then follows using Theorem (2.5). This completes the proof of part 2.